The Mega Division of Green Corporation is an investment center. It has $1,000,000 of operating assets. During 2018, the Mega Division earned operating income of $300,000 on $6,000,000 of sales. Green’s companywide return on investment or desired rate of return is approximately 10%

Please Show Work, and No excel sheets:

  1. What is the ROI?______________________

  1. What is the margin? ____________________

  1. What is the turnover? _____________________

  1. What is the residual income? _______________
Answer #1

ROI = Net operating income / Operating assets

= 300,000 / 1,000,000

= 30%

Margin = Net operating income / Sales

= 300,000 / 6,000,000

= 5%

Turnover = Sales / operating assets

= 6,000,000 / 1,000,000

= 6 times

Residual income = Operating income - required return on investment

= 300,000 - (1,000,000 * 10%)

= 200,000